the width of a rectangle is 8cm shorter than it's length. if the perimeter of the rectangle is 84cm what are it's dimensions
Answer by checkley71(8403)

2X+2(X-8)=84
2X+2X-16=84
4X=84+16
4X=100
X=100/4
X=25 LENGTH
25-8=17 WIDTH
2*25+2*17=84
50+34=84
84=84